# On zero-shot learning in neural state estimation of power distribution systems

### v1

@@ 5.2 Correlation analysis
Looking at the correlation matrix in figure 1, we can spot several relationships:
  • Model complexity is negatively correlated with performance in all use cases except the heterogeneous transfer from PQ to MV. This could already be seen in Ranking model configurations.
  • Feature propagation benefits smaller models and hampers larger ones, as discussed above.
  • Admittance weights generally do not contribute to performance, except in the heterogeneous transfer use case.
  • Performance in OD and TC1 scenarios heavily depends on how well the model can fit on the PQ dataset. However, the opposite is true for the heterogeneous transfer from PQ to MV. This indicates a tradeoff between the accuracy on one dataset and generality.
The negative correlation between model complexity and performance suggests that simpler models may generalize better in most ZSL scenarios. This could be attributed to the fact that complex models tend to overfit the training data, capturing noise rather than the underlying structure, which is detrimental when the model is exposed to new, unseen topologies. Conversely, simpler models with fewer parameters are less prone to overfitting, thereby maintaining better generalization performance.
5.3 Impact of model complexity
Before proceeding to the next analysis, it should be noted that GNNs are models with an inherently low number of trainable parameters. At the lowest end, a single layer of a GCN has only 6 parameters. This is because these models only learn local interactions in the graph, which have few possible configurations.
Figure 2: Performance as a function of model complexity
In figure 2, we plot the performance of models by the composite metric against their trainable parameter counts. We also plot the average for all models of a given complexity: the “Mean” line on the graph.
The most surprising observation here is that there is a clear break point at 72 parameters, starting from which the performance of models becomes much more consistent between use cases as well as each other. Unfortunately, this consistent level of performance is rather low, so shallower models outperform the deeper ones in most use cases.
5.4 Impact of data augmentations
fp adm test_base od tc1 tc2 pq2mv mv2pq
True True 0.526364 0.589545 0.321824 0.324198 0.410957 6.140329
True False 0.529878 0.592244 0.324559 0.324826 0.383784 2.692275
False True 0.521979 0.544307 0.319624 0.320599 0.478989 3.416003
False False 0.522187 0.541419 0.319731 0.321882 0.411902 3.131018
Table 6: Impact of data augmentations
In table 6, we average the value of the composite metric across all model configurations, leaving only the data augmentations as parameters. The results are unsurprising and follow the observations we already made in Correlation analysis. In most use cases, the impact of the augmentations is insignificant, except for heterogeneous transfers, where admittance weights provide a significant improvement.
5.5 Alignment of metrics
To test if the resilience metric we introduced in the Research question section is aligned with MSE, which was used as the loss function for training the models, we computed the Pearson product-moment correlation coefficients between them across all use cases and model configurations. Note that the directions of the metrics are opposite to each other, so we also reverse the sign of the correlation coefficients to avoid confusion.
  • Observability degradation use case: $0.55184$
  • Homogeneous transfer from PQ to TC: $-0.12165$
  • Homogeneous transfer from TC to TC: 00
  • Heterogeneous transfer from PQ to MV: $0.624917$
  • Heterogeneous transfer from MV to PQ: $0.536992$
We observe that the metrics are aligned for observability degradation as well as well as heterogeneous transfers but misaligned for homogeneous transfers. This explains the poor performance in the homogeneous use case and points out to the necessity of developing better metrics and loss functions for NSE.
6 Conclusion
The findings from our research affirm the potential of GNNs in enhancing the adaptability and accuracy of state estimation in power distribution systems. Our experiments demonstrate the GNNs ability to maintain high accuracy when adapted to new topologies, confirming their suitability for dynamic environments where grid configurations frequently change.
The novel conclusions of this paper can be summarized as follows:
@@ 6 Conclusion
    Despite the conventional wisdom in the ML community being "Scale is all you need," here we observe the opposite effect: scaling GNNs up is generally not the best way to improve NSE performance.
@@
    The performance of different model architectures across benchmarks converges at a certain scale but to a low level. Understanding the mechanism of this convergence presents a very interesting research question.
@@
    While the effectiveness of data augmentation techniques such as feature propagation and admittance weights varied significantly across different use cases, they generally improved performance more than scaling the models up did.
@@
    Measuring the performance of NSE by aggregate error alone is not always representative. We attempted to introduce a metric based on real-world use of NSE, but it did not align with aggregated error in the case of homogeneous topology changes. Understanding the reason for that and developing a better metric is another promising research question.
  5. 5.
    ZSL between completely different grid topologies works surprisingly well when transferring from a smaller grid to a bigger one. The reasons for this are unclear, but exploiting this phenomenon could potentially yield improved techniques for training NSE models, such as pretaining on small grids before training on larger ones.
{credits}
6.0.1 Acknowledgments
@@
6.0.2 Availability of data and source code
The source code for this project, along with the datasets, are openly available in the following repository:
https://gitlab.com/transense/nse-tl-paper/tree/ECML

### v2

@@ V-D Correlation analysis
To explore the correlations between hyperparameters and use cases, we compute the Pearson correlation matrix between their associated performance values in Figure 3. Here, we also use the number of trainable parameters (“#params”) instead of layers to compare model complexity more fairly. Note that since lower MSE means better performance, a positive correlation between a hyperparameter and performance is shown as negative and vice versa.
The analysis confirms the conclusions that we made previously: model complexity is detrimental to performance for all use cases except MV2PQ, and our proposed augmentations only marginally affect performance, with feature propagation being most useful in homogeneous scenarios.
VI Limitations and future work
During the evaluation stage of this research, it became evident that MSE alone does not convey enough information to confidently make conclusions about ZSL performance of the models. However, we could not find a better alternative in the literature.
The main problem is that MSE only shows us the instantaneous performance and does not account for the transfer process. An ideal metric $M$ for ZSL and TL experiments would be a differential one that takes into account the magnitude of the change in the underlying grid $\Delta G$ and the performance change $\Delta P$, for example,
$M=\frac{\Delta P}{\Delta G}$
However, an algorithm to compute $\Delta G$ is not trivial to develop. We hope to tackle this problem in our future work.
VII Conclusion
The findings of this paper can be summarized as follows:
@@ VII Conclusion
    GNNs are very robust to homogeneous topology changes in the underlying power grid.
@@
    Some GNNs can perform well in a zero-shot transfer from a larger grid to a smaller one, but not in the other direction.
@@
    Despite the conventional wisdom in the ML community being “Scale is all you need,” scaling GNNs up is not the best way to improve NSE performance. This is explained by the oversmoothing phenomenon [oversmoothing].
@@
    Measuring the performance of NSE by MSE is not always helpful, especially in the context of ZSL. However, we could not find another commonly accepted metric. The development of such a metric appears to be a research gap at the moment.
Acknowledgments
@@
Availability of data and source code
The source code for this project, along with the datasets, is openly available in the following repository:
https://gitlab.com/transense/nse-tl-paper/tree/IARIA
[…]